# Deterministic Dynamic Edge-Colouring

### v1

@@ 4.1 Analysing the Local Condition
Theorem 21.
Let $0<\varepsilon\leq\frac{1}{2}$ and let $G$ be equipped with a $t$-colouring $c$ satisfying Invariant 20. Let $1\leq\eta\leq\tfrac{\varepsilon^{2}}{128\log n}\tfrac{\Delta_{\max}}{t}$. Then it holds that $\tilde{d}\leq(1+\varepsilon)\frac{\Delta_{\max}}{t}$.
Before we give this proof, we first show some helpful lemmas. The first lemma shows that if an edge $e$ has a large surplus, then any endpoint of $e$ has a large surplus in some colour.
Lemma 22.
Suppose Invariant 20 holds. Suppose furthermore that $e=uv$ has $L(uv)=D$. Then there exists a colour $\kappa$ such that $s_{\kappa}(u)\geq D-\eta$.
Proof.
Consider $v$. Since $d(v)\leq\Delta_{\max}$, it follows by the pigeon-hole principle that there exists some colour $\kappa^{\prime}$ for which $d_{\kappa^{\prime}}(v)\leq\frac{\Delta}{t}$. Since we assumed Invariant 20 holds, we have that
@@
Hence, we deduce that $D-\eta\leq s_{\kappa^{\prime}}(u)$, and we are done by setting $\kappa=\kappa^{\prime}$. ∎
@@
Lemma 23.
Suppose Invariant 20 holds, and let $\mu\in\mathbb{Z}$ be a parameter. Then there exists two colours $\kappa_{x}$ and $\kappa_{y}$, as well as two family of sets $(X_{i})_{i=1}^{\mu}$ and $(Y_{i})_{i=1}^{\mu}$ such that for all $i$, we have $X_{i},Y_{i}\subset V(G)$ and that the following two conditions hold:
$\forall u\in X_{i}:s_{\kappa_{x}}(u)-s_{\kappa_{y}}(u)$ $\geq\tilde{L}-(1+2(i-1))\eta$ (1)
$\forall v\in Y_{i}:s_{\kappa_{y}}(v)-s_{\kappa_{x}}(v)$ $\geq\tilde{L}-2i\eta$ (2)
Proof.
We first determine $\kappa_{x}$ and $\kappa_{y}$. Then we construct the sets inductively. Pick an arbitrary edge $uv$ achieving $L(uv)=\max L(e)=\tilde{L}$. Now we may apply Lemma 22 to see that there exists a colour $\kappa_{x}$ such that $d_{\kappa_{x}}(u)\geq\frac{\Delta_{\max}}{t}+\tilde{L}-\eta$. By the pigeon-hole principle there is some colour $\kappa_{y}$ incident to $u$ such that $d_{\kappa_{y}}(u)\leq\frac{\Delta_{\max}}{t}$.
With $\kappa_{x}$ and $\kappa_{y}$ defined, we construct the family of sets inductively. Initially, we set $X_{1}=\{u\}$ and $Y_{1}=N_{G[\kappa_{x}]}(u)-\{u\}$. We confirm that the required conditions hold: $(1)$ is clear from our choice of $\kappa_{x}$ and $\kappa_{y}$. We get $(2)$ by observing that by Invariant 20, we find that for any $v\in Y_{1}$, we have:
@@
For the inductive step, we now set $X_{2}=N_{G[\kappa_{y}]}(Y_{1})-Y_{1}$, and $Y_{2}=N_{G[\kappa_{x}]}(X_{2})-X_{2}$. Again we confirm the two conditions:
  • $(1)$: Consider $u\in X_{i}$. We have $\forall v\in Y_{i-1}$ that $s_{\kappa_{y}}(v)-s_{\kappa_{x}}(v)\geq\tilde{L}-2(i-1)\eta$ by the induction hypothesis. By applying Invariant 20, we find that
@@
    $s_{\kappa_{x}}(u)-s_{\kappa_{y}}(u)$ $\geq s_{\kappa_{x}}(v)-s_{\kappa_{y}}(v)-\eta$
@@
  • $(2)$ is symmetric: consider any $v\in Y_{i}$. We have $\forall u\in X_{i}$ that $s_{\kappa_{x}}(u)-s_{\kappa_{y}}(u)\geq\tilde{L}-(1+2(i-1))\eta$ by induction. By applying Invariant 20, we find that
@@
We are now ready to show Theorem 21:
Proof of Theorem 21.
Suppose for contradiction that $\tilde{d}>(1+\varepsilon)\frac{\Delta}{t}$. We will show that this implies the existence of a family of bipartite subgraphs $\{\mathbb{B}_{j}\}_{j=1}^{10\log(n)\varepsilon^{-1}}=\{(P_{j},Q_{j}))\}_{j=1}^{10\log(n)\varepsilon^{-1}}$ of $G$ satisfying that $|Q_{j}|\geq(1+\frac{\varepsilon}{4})^{j}$.
@@
where we used the standard in-equality $1-x\leq e^{-x}$ which holds for all real $x$. In particular, it follows that for all $x\leq 1/2$, we have that $e^{x}\leq\frac{1}{1-x}\leq 1+\frac{x}{1-x}\leq 1+2x$. This contradicts the fact that we also must have $\mathbb{B}_{j}\subset G$ and so $|V(\mathbb{B}_{j})|\leq|V(G)|=n$.
In order to construct $\{\mathbb{B}_{j}\}_{j=1}^{10\log(n)\varepsilon^{-1}}$, we invoke Lemma 23 with $\mu=\lceil 10\log(n)\varepsilon^{-1}\rceil$ to get sets $(X_{i})_{i=1}^{\mu}$ and $(Y_{i})_{i=1}^{\mu}$ and colours $\kappa_{x}$ and $\kappa_{y}$ satisfying the condition of the lemma. Then for appropriate values of $i$, we set
@@
Instead of showing that $Q_{2}$ is also sufficiently large, we proceed to the induction step directly. We will first show that if $Q_{2i-1}\geq(1+\frac{\varepsilon}{4})^{2i-1}$, then $Q_{2i}\geq(1+\frac{\varepsilon}{4})^{2i}$. Then we show that if $Q_{2i}\geq(1+\frac{\varepsilon}{4})^{i}$ , then $Q_{2i+1}\geq(1+\frac{\varepsilon}{4})^{2i+1}$. In this way, we only need to show the base case $|Q_{1}|=|Y_{1}|\geq\tilde{L}-\eta\geq 2$, as we have already done.
@@
$((1+2i)\eta+\frac{\Delta}{t})|X_{i+1}|\geq\sum\limits_{u\in X_{i+1}}d_{\kappa_{y}}(u)\geq\sum\limits_{v\in Y_{i}}d_{\kappa_{y}}(v)\geq(\tilde{L}-2i\eta+\frac{\Delta}{t})|Y_{i}|$
Since $\tilde{d}\leq\frac{\Delta}{t}+\tilde{L}$ by definition and since $\eta\frac{22\log(n)}{\varepsilon}\leq\frac{\varepsilon\Delta}{4t}$ by choice of $\eta$, we have that
$\tilde{L}+\frac{\Delta}{t}-\eta\frac{22\log(n)}{\varepsilon}\geq\tilde{d}-\eta\frac{22\log(n)}{\varepsilon}\geq(1+\frac{3\varepsilon}{4})\frac{\Delta}{t}$
@@
$\frac{\Delta}{t}+\eta\frac{22\log(n)}{\varepsilon}\leq(1+\frac{\varepsilon}{4})\frac{\Delta}{t}$
@@
$(1+\frac{\varepsilon}{4})\frac{\Delta}{t}|X_{i+1}|\geq(1+\frac{3\varepsilon}{4})\frac{\Delta}{t}|Y_{i}|$
@@
This implies that every vertex $v\in Y_{i}$ has $s_{\kappa_{x}}(v)\leq 2(i+1)\eta$ and that every vertex $u\in X_{i+1}$ has $s_{\kappa_{x}}(u)\geq\tilde{L}-(1+2i)\eta$. Since the graph is bipartite, we have that
@@
$(2(i+1)\eta+\frac{\Delta}{t})|Y_{i+1}|\geq\sum\limits_{u\in X_{i+1}}d_{\kappa_{y}}(u)\geq\sum\limits_{v\in Y_{i}}d_{\kappa_{y}}(v)\geq(\tilde{L}-(1+2i)\eta+\frac{\Delta}{t})|X_{i+1}|$
Since $\tilde{d}\leq\frac{\Delta}{t}+\tilde{L}$ by definition and since $\eta\frac{22\log(n)}{\varepsilon}\leq\frac{\varepsilon\Delta}{4t}$ by choice of $\eta$, we have similarly to before that
$\tilde{L}+\frac{\Delta}{t}-\eta\frac{22\log(n)}{\varepsilon}\geq\tilde{d}-\eta\frac{22\log(n)}{\varepsilon}\geq(1+\frac{3\varepsilon}{4})\frac{\Delta}{t}$
@@
$\frac{\Delta}{t}+\eta\frac{22\log(n)}{\varepsilon}\leq(1+\frac{\varepsilon}{4})\frac{\Delta}{t}$
@@
$(1+\frac{\varepsilon}{4})\frac{\Delta}{t}|Y_{i+1}|\geq(1+\frac{3\varepsilon}{4})\frac{\Delta}{t}|X_{i+1}|$
@@
since $Q_{2i+1}=Y_{i+1}$ and $Q_{2i}=X_{i+1}$, applying the induction hypothesis thus yields the second case. ∎
Note that Theorem 21 shows that any colouring satisfying the local condition is a $t$-degree splitting colouring.
4.2 Dynamically Maintaining Invariant 20
@@
    $\tilde{s}_{c(wx)}(w)+\tilde{s}_{c(wx)}(x)-(\tilde{s}_{\kappa}(w)+\tilde{s}_{\kappa}(x))$
@@
The algorithm now works as follows (see below for pseudo-code). Whenever an edge $e=uv$ is inserted, we calculate the surplus $s_{\kappa}(u)+s_{\kappa}(v)$ for all choices of $\kappa$, and assign $e$ an arbitrary colour $\kappa^{\prime}$ with minimum surplus. We also initialize all of $e$’s data structures, and update its contributions to $u$ and $v$’s data structures. Then, we insert $e$ into the linked-lists $LL(u,\kappa^{\prime})$ and $LL(v,\kappa^{\prime})$, ensuring that $e$ is inserted just before position $\iota(u,\kappa^{\prime})$ in $LL(u,\kappa^{\prime})$ resp.  $\iota(v,\kappa^{\prime})$ in $LL(v,\kappa^{\prime})$. Then we update data structures for other edges $f$ around $w\in\{u,v\}$ in a round-robin fashion: starting with the edge stored at position $\iota(w,\kappa^{\prime})$ in $LL(w,\kappa^{\prime})$, we recompute the contents of the next $\lceil 500\tfrac{\Delta_{\max}}{\eta}\rceil$ edges in $LL(w,\kappa^{\prime})$ and update $\iota(w,\kappa^{\prime})=\iota(w,\kappa^{\prime})+\lceil 500\tfrac{\Delta_{\max}}{\eta}\rceil\mod d_{\kappa^{\prime}}(w)$. Then, we check for all colours if any violations have arisen. If so, we push the endpoint and the colour with too large a surplus difference, unto a queue $Q$. After this, we process $Q$ one pair at the time by recolouring with a similar procedure as above, possibly pushing new violating pairs unto $Q$. Note that to get the edge to recolour, we extract the observed minimum and check if its surplus difference is still large enough. If yes, we process it by recolouring it, if no, we conclude our processing of this element of $Q$. Deletions are symmetric. Here, we have to remember to remove $e$ from any doubly-linked-lists containing it, and to re-adjust the relevant $\iota(\cdot{},\cdot{})$ by decrementing them by $1$, if $e$’s position in $LL(\cdot{},\cdot{})$ was before the element pointed to by $\iota(\cdot{},\cdot{})$.
@@
Note that this value can be found by querying $MSD_{\kappa}(w)$ and returning the minimum.

### v2

@@ 4.1 Analysing the Local Condition
Theorem 4.3.
Let $0<\varepsilon\leq\frac{1}{2}$ and let $G$ be equipped with a $t$-colouring $c$ satisfying Invariant 4.2. Let $1\leq\eta\leq\tfrac{\varepsilon^{2}}{128\log n}\tfrac{\Delta_{\max}}{t}$. Then it holds that $\tilde{d}\leq(1+\varepsilon)\frac{\Delta_{\max}}{t}$.
Before we show this theorem, we first show some helpful lemmas. The first lemma shows that if an edge $e$ has a large surplus, then any endpoint of $e$ has a large surplus in some colour.
Lemma 4.4.
Suppose Invariant 4.2 holds. Suppose furthermore that $e=uv$ has $L(uv)=D$. Then there exists a colour $\kappa$ such that $s_{\kappa}(u)\geq D-\eta$.
Proof 5.
Consider $v$. Since $d(v)\leq\Delta_{\max}$, it follows by the pigeon-hole principle that there exists some colour $\kappa^{\prime}$ for which $d_{\kappa^{\prime}}(v)\leq\frac{\Delta_{\max}}{t}$. Since we assumed Invariant 4.2 holds, we have that
@@
Hence, we deduce that $D-\eta\leq s_{\kappa^{\prime}}(u)$, and we are done by setting $\kappa=\kappa^{\prime}$.
@@
Lemma 4.5.
Suppose Invariant 4.2 holds, and let $\mu\in\mathbb{Z}$ be a parameter. Then there exists two colours $\kappa_{x}$ and $\kappa_{y}$, as well as two family of sets $(X_{i})_{i=1}^{\mu}$ and $(Y_{i})_{i=1}^{\mu}$ such that for all $i$, we have $X_{i},Y_{i}\subset V(G)$ and that the following two conditions hold:
(4.1) $\forall u\in X_{i}:s_{\kappa_{x}}(u)-s_{\kappa_{y}}(u)$ $\geq\tilde{L}-(1+2(i-1))\eta$
(4.2) $\forall v\in Y_{i}:s_{\kappa_{y}}(v)-s_{\kappa_{x}}(v)$ $\geq\tilde{L}-2i\eta$
Proof 6.
We first determine $\kappa_{x}$ and $\kappa_{y}$. Then we construct the sets inductively. Pick an arbitrary edge $uv$ achieving $L(uv)=\max L(e)=\tilde{L}$. Now we may apply Lemma 4.4 to see that there exists a colour $\kappa_{x}$ such that $d_{\kappa_{x}}(u)\geq\frac{\Delta_{\max}}{t}+\tilde{L}-\eta$. By the pigeon-hole principle there is some colour $\kappa_{y}$ incident to $u$ such that $d_{\kappa_{y}}(u)\leq\frac{\Delta_{\max}}{t}$.
With $\kappa_{x}$ and $\kappa_{y}$ defined, we construct the family of sets inductively. Initially, we set $X_{1}=\{u\}$ and $Y_{1}=N_{G[\kappa_{x}]}(u)-\{u\}$. We confirm that the required conditions hold: $(1)$ is clear from our choice of $\kappa_{x}$ and $\kappa_{y}$. We get $(2)$ by observing that by Invariant 4.2, we find that for any $v\in Y_{1}$, we have:
@@
For the inductive step, we now set $X_{i}=N_{G[\kappa_{y}]}(Y_{i-1})-Y_{i-1}$, and $Y_{i}=N_{G[\kappa_{x}]}(X_{i})-X_{i}$. Again we confirm the two conditions:
  • $(1)$: Consider $u\in X_{i}$. We have $\forall v\in Y_{i-1}$ that $s_{\kappa_{y}}(v)-s_{\kappa_{x}}(v)\geq\tilde{L}-2(i-1)\eta$ by the induction hypothesis. By applying Invariant 4.2, we find that
@@
    $s_{\kappa_{x}}(u)-s_{\kappa_{y}}(u)$ $\geq s_{\kappa_{y}}(v)-s_{\kappa_{x}}(v)-\eta$
@@
  • $(2)$ is symmetric: consider any $v\in Y_{i}$. We have $\forall u\in X_{i}$ that $s_{\kappa_{x}}(u)-s_{\kappa_{y}}(u)\geq\tilde{L}-(1+2(i-1))\eta$ by induction. By applying Invariant 4.2, we find that
@@
We are now ready to show Theorem 4.3:
Proof 7 (Proof of Theorem 4.3).
Suppose for contradiction that $\tilde{d}>(1+\varepsilon)\frac{\Delta_{\max}}{t}$. We will show that this implies the existence of a family of bipartite subgraphs $\{\mathbb{B}_{j}\}_{j=1}^{10\log(n)\varepsilon^{-1}}=\{(P_{j},Q_{j})\}_{j=1}^{10\log(n)\varepsilon^{-1}}$ of $G$ satisfying that $|Q_{j}|\geq(1+\frac{\varepsilon}{4})^{j}$.
@@
where we used the standard in-equality $1-x\leq e^{-x}$ which holds for all real $x$. In particular, it follows that for all $0\leq x\leq 1/2$, we have that $e^{x}\leq\frac{1}{1-x}\leq 1+\frac{x}{1-x}\leq 1+2x$. This contradicts the fact that we also must have $\mathbb{B}_{j}\subset G$ and so $|V(\mathbb{B}_{j})|\leq|V(G)|=n$.
In order to construct $\{\mathbb{B}_{j}\}_{j=1}^{10\log(n)\varepsilon^{-1}}$, we invoke Lemma 4.5 with $\mu=\lceil 10\log(n)\varepsilon^{-1}\rceil$ to get sets $(X_{i})_{i=1}^{\mu}$ and $(Y_{i})_{i=1}^{\mu}$ and colours $\kappa_{x}$ and $\kappa_{y}$ satisfying the condition of the lemma. Then for appropriate values of $i$, we set
@@
Instead of showing that $Q_{2}$ is also sufficiently large, we proceed to the induction step directly. We will first show that if $Q_{2i-1}\geq(1+\frac{\varepsilon}{4})^{2i-1}$, then $Q_{2i}\geq(1+\frac{\varepsilon}{4})^{2i}$. Then we show that if $Q_{2i}\geq(1+\frac{\varepsilon}{4})^{2i}$ , then $Q_{2i+1}\geq(1+\frac{\varepsilon}{4})^{2i+1}$. In this way, we only need to show the base case $|Q_{1}|=|Y_{1}|\geq\tilde{L}-\eta\geq 2$, as we have already done.
@@
$((1+2i)\eta+\frac{\Delta_{\max}}{t})|X_{i+1}|\geq\sum\limits_{u\in X_{i+1}}d_{\kappa_{y}}(u)\geq\sum\limits_{v\in Y_{i}}d_{\kappa_{y}}(v)\geq(\tilde{L}-2i\eta+\frac{\Delta_{\max}}{t})|Y_{i}|$
Since $\tilde{d}\leq\frac{\Delta_{\max}}{t}+\tilde{L}$ by definition and since $\eta\frac{22\log(n)}{\varepsilon}\leq\frac{\varepsilon\Delta_{\max}}{4t}$ by choice of $\eta$, we have that
$\tilde{L}+\frac{\Delta_{\max}}{t}-\eta\frac{22\log(n)}{\varepsilon}\geq\tilde{d}-\eta\frac{22\log(n)}{\varepsilon}\geq(1+\frac{3\varepsilon}{4})\frac{\Delta_{\max}}{t}$
@@
$\frac{\Delta_{\max}}{t}+\eta\frac{22\log(n)}{\varepsilon}\leq(1+\frac{\varepsilon}{4})\frac{\Delta_{\max}}{t}$
@@
$(1+\frac{\varepsilon}{4})\frac{\Delta_{\max}}{t}|X_{i+1}|\geq(1+\frac{3\varepsilon}{4})\frac{\Delta_{\max}}{t}|Y_{i}|$
@@
This implies that every vertex $v\in Y_{i+1}$ has $s_{\kappa_{x}}(v)\leq 2(i+1)\eta$ and that every vertex $u\in X_{i+1}$ has $s_{\kappa_{x}}(u)\geq\tilde{L}-(1+2i)\eta$. Since the graph is bipartite, we have that
@@
$(2(i+1)\eta+\frac{\Delta_{\max}}{t})|Y_{i+1}|\geq\sum\limits_{u\in X_{i+1}}d_{\kappa_{y}}(u)\geq\sum\limits_{v\in Y_{i}}d_{\kappa_{y}}(v)\geq(\tilde{L}-(1+2i)\eta+\frac{\Delta_{\max}}{t})|X_{i+1}|$
Since $\tilde{d}\leq\frac{\Delta_{\max}}{t}+\tilde{L}$ by definition and since $\eta\frac{22\log(n)}{\varepsilon}\leq\frac{\varepsilon\Delta_{\max}}{4t}$ by choice of $\eta$, we have similarly to before that
$\tilde{L}+\frac{\Delta_{\max}}{t}-\eta\frac{22\log(n)}{\varepsilon}\geq\tilde{d}-\eta\frac{22\log(n)}{\varepsilon}\geq(1+\frac{3\varepsilon}{4})\frac{\Delta_{\max}}{t}$
@@
$\frac{\Delta_{\max}}{t}+\eta\frac{22\log(n)}{\varepsilon}\leq(1+\frac{\varepsilon}{4})\frac{\Delta_{\max}}{t}$
@@
$(1+\frac{\varepsilon}{4})\frac{\Delta_{\max}}{t}|Y_{i+1}|\geq(1+\frac{3\varepsilon}{4})\frac{\Delta_{\max}}{t}|X_{i+1}|$
@@
since $Q_{2i+1}=Y_{i+1}$ and $Q_{2i}=X_{i+1}$, applying the induction hypothesis thus yields the second case.
Note that Theorem 4.3 shows that any colouring satisfying the local condition is a $t$-degree splitting colouring.
4.2 Dynamically Maintaining Invariant 4.2
@@
    $\tilde{s}_{c(ux)}(u)+\tilde{s}_{c(ux)}(x)-(\tilde{s}_{\kappa}(u)+\tilde{s}_{\kappa}(x))$
@@
The algorithm now works as follows (see below for pseudo-code). Whenever an edge $e=uv$ is inserted, we calculate the surplus $s_{\kappa}(u)+s_{\kappa}(v)$ for all choices of $\kappa$, and assign $e$ an arbitrary colour $\kappa^{\prime}$ with minimum surplus. We also initialize all of $e$’s data structures, and update its contributions to $u$ and $v$’s data structures. Then, we insert $e$ into the linked-lists $LL(u,\kappa^{\prime})$ and $LL(v,\kappa^{\prime})$, ensuring that $e$ is inserted just before position $\iota(u,\kappa^{\prime})$ in $LL(u,\kappa^{\prime})$ resp. $\iota(v,\kappa^{\prime})$ in $LL(v,\kappa^{\prime})$. Then we update data structures for other edges $f$ around $w\in\{u,v\}$ in a round-robin fashion: starting with the edge stored at position $\iota(w,\kappa)$ in $LL(w,\kappa)$, we recompute the contents of the next $\lceil 500\tfrac{\Delta_{\max}}{\eta}\rceil$ edges in $LL(w,\kappa)$ and update $\iota(w,\kappa)=\iota(w,\kappa)+\lceil 500\tfrac{\Delta_{\max}}{\eta}\rceil\mod d_{\kappa}(w)$. Note that this is done for all choices of colours $\kappa$.
Then, we check for all colours if any violations have arisen. If so, we push the endpoint and the colour with too large a surplus difference, unto a queue $Q$. After this, we process $Q$ one pair at the time by recolouring with a similar procedure as above, possibly pushing new violating pairs unto $Q$. Note that to get an edge to recolour, we extract the observed maximum and check if its surplus difference is still large enough. If yes, we process it by recolouring it, if no, we conclude our processing of this element of $Q$. Deletions are symmetric in the sense that we simply remove the edge and update the affected data structures, before we update in a round-robin fashion and fix violations exactly as before. Here, we have to remember to remove $e$ from any doubly-linked-lists containing it, and to re-adjust the relevant $\iota(\cdot{},\cdot{})$ by decrementing them by $1$, if $e$’s position in $LL(\cdot{},\cdot{})$ was before the element pointed to by $\iota(\cdot{},\cdot{})$.
@@
Note that this value can be found by querying $MSD_{\kappa}(w)$ and returning the maximum.
$\texttt{add}(uv)$ to $G$
$\texttt{Initialize-DS}(uv)$
$Q\leftarrow\texttt{ReColour}(uv,\emptyset)$
while $Q\neq\emptyset$ do
  $(w,\kappa)\leftarrow\texttt{pop}(Q)$
  if $\Delta\tilde{s}_{\kappa}(w)>\frac{\eta}{2}$ then
   $e\leftarrow$ edge maximising the above expression.
   $Q\leftarrow\texttt{ReColour}(e,Q)$
  end if
end while
Algorithm 1 Insert(uv)(uv)
$\texttt{Remove}(uv)$ from $G$.
 Update uvuv’s influence on relevant $SP(\cdot{})$$LL(\cdot{},\cdot{})$, $\iota(\cdot{},\cdot{})$, $MSD_{\cdot{}}(\cdot{})$
$\texttt{Round-Robin}(uv)$
for all $\kappa\in[t]$ and $w\in\{u,v\}$ do
  if $\Delta\tilde{s}_{\kappa}(w)>\frac{\eta}{2}$ then
   $\texttt{Push}((w,\kappa))$ to $Q$
  end if
end for
while $Q\neq\emptyset$ do
  $(w,\kappa)\leftarrow\texttt{pop}(Q)$
  if $\Delta\tilde{s}(w)>\frac{\eta}{2}$ then
   $e\leftarrow$ edge maximising the above expression.
   $Q\leftarrow\texttt{ReColour}(e,Q)$
  end if
end while
@@
if $c(uv)=\emptyset$ then
  $c(uv)\leftarrow\texttt{ExtractMinColour}(T^{s}_{uv})$
  Update uvuv’s influence on relevant $SP(\cdot{})$, $LL(\cdot{},\cdot{})$, $\iota(\cdot{},\cdot{})$, $MSD_{\cdot{}}(\cdot{})$
  $\texttt{Update-DS}(uv)$
else
  $c(uv)\leftarrow\texttt{ExtractMinColour}(T^{s}_{uv})$
  Update uvuv’s influence on relevant $SP(\cdot{})$, $LL(\cdot{},\cdot{})$, $\iota(\cdot{},\cdot{})$, $MSD_{\cdot{}}(\cdot{})$
  $\texttt{Update-DS}(uv)$
end if
$\texttt{Round-Robin}(uv)$
for all $\kappa\in[t]$ and $w\in\{u,v\}$ do
  if $\Delta\tilde{s}_{\kappa}(w)>\frac{\eta}{2}$ then
   $\texttt{Push}((w,\kappa))$ to $Q$
  end if
end for
Return $Q$
Algorithm 3 ReColour(uv,Q)(uv,Q)
for all $\kappa\in[t]$ and $w\in\{u,v\}$ do
  for all $j\in[\lceil\tfrac{500\Delta_{\max}}{\eta}\rceil]$ do
   $e\leftarrow LL(w,\kappa)[\iota(w,\kappa)+(j-1)\mod d_{\kappa}(w)]$
   $\texttt{Update-DS}(e)$
  end for
  $\iota(w,\kappa)\leftarrow\iota(w,\kappa)+\lceil\tfrac{500\Delta_{\max}}{\eta}\rceil\mod d_{\kappa}(w)$
  Update pointer associated with $\iota(w,\kappa)$
end for
Algorithm 4 Round-Robin(uv)(uv)
if $d(w)=1$ for $w\in\{u,v\}$ then
  Initialize DS for $w$
end if
 Initialize BBST $T^{s}_{uv}$.
for all $\kappa\in[t]$ do
@@
end for
Algorithm 5 Initialize-DS(uv)(uv)
 Update $SP(u)$ and $SP(v)$
for all $\kappa\in[t]$ do
  Update ($s_{\kappa}(u)+s_{\kappa}(v)$) in $T^{s}_{uv}$
end for
for all $\kappa\in[t]$ and $w\in\{u,v\}$ do
  Update uvuv’s contribution to $MSD_{\kappa}(w)$
end for
Algorithm 6 Update-DS(uv)(uv)